OpenStudy (alexis):

What is the length of the missing side of a right triangle with 10 units as the hypotenuse and 6 units as one of the sides? A. 64 units B. 4 units C. 8 units D. 16 units

OpenStudy (jwhorne411):

You can use the equation \[c^2=a^2+b^2\] to solve this. where c = 10, a = 6, and solve for b. \[b=\sqrt{c^2-a^2}\]. Therefore \[b=\sqrt{100-36}=8\] . The correct choice is C. 8 units.